nLet (a

Let (an) n belongs to N be any sequence. Show that limn--->∞ an= L iff for every ε > 0 there exists some N belongs to N

such that n ≥ N implies an belongs to Nε (L)


1
Expert's answer
2021-03-26T16:41:20-0400

Solution:

Let "\left\{a_{n}\right\}" be a sequence with positive terms such that "\lim _{n \rightarrow \infty} a_{n}=L>0" .

 Let "x" be a real number.

To prove that "\lim _{n \rightarrow \infty} a_{n}^{x}=L^{x}" .

Proof: Let "\epsilon>0". Note that "L<\left(L^{x}+\epsilon\right)^{1 / x}" and "L>\left(L^{x}-\epsilon\right)^{1 / x}" .

Since "\lim _{n \rightarrow \infty} a_{n}=L" , there exists some N such that "n \geq N \implies a_{n}<\left(L^{x}+\epsilon\right)^{1 / x}"

and "a_{n}>\left(L^{x}-\epsilon\right)^{1 / x}"

Hence, for "n \geq N" we have "a_{n}^{x}<L^{x}+\epsilon"

and "a_{n}^{x}>L^{x}-\epsilon" . This shows "\lim _{n \rightarrow \infty} a_{n}^{x}=L^{x}" .

Hence, proved.